Differentiate root x^ x ^x^ x^ x^ x^ x ^x ^x....................upto infinity w.r.t. x

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
we can write it as y = sqrt(x *sqrt (y))
 
y^2 = x * sqrt (y)
 
2 y dy/dx = x / (2 *sqrt(y)) * dy/dx + sqrt (y)
 
Hope it helps